OpenStudy (anonymous):

x+4/x^2-4 - 15/x-2 Subtracting expressions.

jhonyy9 (jhonyy9):

x^2 -4=(x-2)(x+2)

OpenStudy (anonymous):

how?

jhonyy9 (jhonyy9):

so (x+4-15(x+2))/(x-2)(x+2) =(x+4-15x-30)/(x-2)(x+2) =(-14x-26)/(x-2)(x+2) = =-2(7x+13)/(x-2)(x+2)

jhonyy9 (jhonyy9):

x^2 -4 =(x-2)(x+2) -us formula a^2 -b^2 =(a-b)(a+b)

jhonyy9 (jhonyy9):

ok ?

OpenStudy (mertsj):